Consider the equation "\frac{x}{2+i}+\frac{iy}{2-i}=\frac{3}{1-2i}" , where x and y are real numbers. Which of the following are correct?


  1. x=−4 and y =-5
  2. x=5 and y=4
  3. x and y satisfy the pair of equations  2xy=3 and -x+2y=6
  4. x=4 and y=5
  5. x and y satisfy the pair of equations 2x+y=3 and x-2y=-6
1
Expert's answer
2021-10-18T15:14:55-0400

1)

L.H.S="\frac{x}{2+i}+\frac{iy}{2-i}"

put x=-4 and y=-5.

"\frac{-4}{2+i}+\frac{-i5}{2-i}\\
\frac{-4(2-i)-5i(2+i)}{(2+i)(2-i)}\\
\frac{-4(2-i)-5i(2+i)}{2^2-i^2}\\
\frac{-3-6i}{5}\\"

Not equal to RHS.

2)

put x=5 and y=4.

"LHS=\frac{5}{2+i}+\frac{i4}{2-i}\\
\frac{5(2-i)+4i(2+i)}{(2+i)(2-i)}\\
\frac{5(2-i)+4i(2+i)}{2^2-i^2}\\
\frac{6+3i}{5}\\"

Not equal to RHS.

3)

solving equation

2x-y=3 and -x+2y=-6, we get x=4, y=5.

"LhS=\frac{4}{2+i}+\frac{i5}{2-i}\\
\frac{4(2-i)+5i(2+i)}{(2+i)(2-i)}\\
\frac{4(2-i)+5i(2+i)}{2^2-i^2}\\
\frac{3+6i}{5}\\"

Not equal to RHS.

4)

putx=4,y=5.

"LhS=\frac{4}{2+i}+\frac{i5}{2-i}\\
\frac{4(2-i)+5i(2+i)}{(2+i)(2-i)}\\
\frac{4(2-i)+5i(2+i)}{2^2-i^2}\\
\frac{3+6i}{5}\\"

Not equal to RHS.

5)

Solving equations 2x+y=3 and x-2y=-6, we get x=0,y=3.

"LhS=\frac{0}{2+i}+\frac{3i}{2-i}\\
\frac{3i}{2-i}\\
\text{By rationalise, we get}\\
\frac{-3+6i}{5}\\"

Not equal to RHS.


No options are correct.